OpenStudy (anonymous):

anyone know of any suggestions (tricks) on remembering trig for calc (i.e. unit circle, identities, etc)

OpenStudy (jdoe0001):

yes, use a cheat sheet :P

OpenStudy (jdoe0001):

heck I have several

OpenStudy (anonymous):

Remember the following identity: \[\sin^2 + \cos^2 = 1\]This is the most important trig identity. You can find most other identities by manipulation this one. For example, you could divide both sides of the equation by cos^2 to get this identity: \[\tan^2 + 1 = \sec^2\]and so on

OpenStudy (anonymous):

nice, thank you

OpenStudy (anonymous):

Here is the cheat sheet I use..